you entered a eurodollar futures position earlier today as the lender (long position) at a price of 94.98. the price has closed to 95.05. How much has your position gained or lost?

A. $700

B. $175

C. $-175

D. $-350

E. $-700
Answer #1

gain on the position=(selling price-purchase price)*100*25=(95.05-94.98)*100*25=175